Most people perceive light having a wavelength between 630nm and 700nm as red and light with a wavelength between 400nm and 440nm as violet.

Part A

Calculate the approximate frequency range for violet light.

Answer in the order indicated. Separate your answers using comma.

Part B

Calculate the approximate frequency range for red light.

Answer in the order indicated. Separate your answers using comma.
Answer #1

Part A

velocity = frequency * wavelength

velocity of light \(=3 \times 10^{8} \mathrm{~m} / \mathrm{s}\) (approx)

wavelength of violet light \(=400 \times 10^{-9} \mathrm{~m}\) to \(440 \times 10^{-9} \mathrm{~m} \quad\) (given)

frequency Range \(=\frac{3 \times 10^{8}}{440 \times 10^{-9}}\) hz to \(\frac{3 \times 10^{8}}{400 \times 10^{-9}} \mathrm{hz}\)

hence frequency Range of violet=.68\times10 \(^{15} \mathrm{hz}, 75 \times 10^{15} \mathrm{hz}\) ans

Part B

wavelength of red light \(=630 \times 10^{-9} \mathrm{~m}\) to \(700 \times 10^{-9} \mathrm{~m} \quad\) (given)

Frequency Range of red light \(=\frac{3 \times 10^{8}}{700 \times 10^{-9}} \mathrm{hz}, \frac{3 \times 10^{8}}{630 \times 10^{-9}} \mathrm{hz}\)